Given


g-10 m/s

media%2F4a3%2F4a338bb4-c0ba-4593-adcf-0c

media%2Fc56%2Fc5620dbd-4d68-41a2-9b92-4e
Problem

media%2Fbe6%2Fbe66eff6-f785-4222-9158-d8
Without using a calculator

A) draw and label free body diagrams for both masses after the rod is removed

B) if the coefficient of static friction us= 0.5, find the minimum pushing force required to keep the 1-kg mass from sliding down the face of the cart
Answer #1

A)

1Kq FeD vo nara

B) To prevent the mass from sliding down, force must be balanced in the vertical direction.

fr = mg

where fr is the frictional force on the mass and mg is the gravitational force.

now

fr = \mu N , N is the normal reaction force.

\therefore \mu N = mg

\therefore N = mg/\mu

Now we write the force equation of the cart

F - N = Ma

as we need to find the minimum force required to prevent the sliding of the mass, the acceleration of the cart should be 0.

\therefore F = N

\therefore F =mg/\mu

Put the values given m = 1kg. g = 10 m/s2\mu = 0.5

\therefore F =1*10/.5

\therefore F = 20 N